When Iron 2 hydroxide is mixed with phosphoric acid, iron 2 phosphate and water are formed. Write the balance equation foe this reaction. If 3.20g of iron 2 hydroxide I'd treated with 2.50g of phosphoric acid, what is the limiting reagent

1
Expert's answer
2022-05-19T03:48:04-0400

Solution:

The molar mass of Fe(OH)2 is 89.86 g/mol

The molar mass of H3PO4 is 97.994 g/mol


Calculate moles of each reactant:

(3.20 g Fe(OH)2) × (1 mol Fe(OH)2 / 89.86 g Fe(OH)2) = 0.0356 mol Fe(OH)2

(2.50 g H3PO4) × (1 mol H3PO4 / 97.994 g H3PO4) = 0.0255 mol H3PO4


Balanced chemical equation:

3Fe(OH)2 + 2H3PO4 → Fe3(PO4)2 + 6H2O

According to stoichiometry:

3 mol of Fe(OH)2 react with 2 mol of H3PO4

Therefore, 0.0356 mol of Fe(OH)2 react with:

(0.0356 mol Fe(OH)2) × (2 mol H3PO4 / 3 mol Fe(OH)2) = 0.0237 mol H3PO4

However, initially there is 0.0255 mol of H3PO4 (according to the task).

Thus, Fe(OH)2 acts as limiting reagent and H3PO4 is excess reagent.


Answer:

Balance equation for the reaction: 3Fe(OH)2 + 2H3PO4 → Fe3(PO4)2 + 6H2O

Iron(II) hydroxide (Fe(OH)2) is the limiting reagent